OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 6 x = –y + 2

OpenStudy (anonymous):

there is There is no solutions. correct?

OpenStudy (anonymous):

A. The solution is (4, 2). B. There is no solution. C. The solution is (2, 4). D. There are an infinite number of solutions.

OpenStudy (anonymous):

yeah, i think you are right, no solutions :)

OpenStudy (anonymous):

thank you! :)

OpenStudy (phi):

if you re-write the second equation x + y = 6 x = –y + 2 to x + y = 6 x +y = 2 you are looking for two numbers that when you add them, you get 6 and *at the same time* add up to 2 there are no numbers that can do that. so there is no solution.

OpenStudy (anonymous):

Thank you @phi

OpenStudy (phi):

or, if you rewrite both equations like this x + y = 6 x = –y + 2 to y = -1x +6 y = -1x + 2 you have the equation of two lines that have the same slope. they are parallel and *never meet* the solution of these two equations is where the lines intersect... but in this case, they never intersect... so no solution.